# Connecting Large Language Models with Evolutionary Algorithms Yields Powerful Prompt Optimizers

### v2

@@ C.4 Analysis of Prompt
Please perform Sentiment Classification task. Given the sentence, assign a sentiment label from [’negative’, ’positive’]. Return label only without any other text.
93.68
Natural Instruction
@@
In this task, you are given sentences from movie reviews. The task is to classify a sentence as "great" if the sentiment of the sentence is positive or as "terrible" if the sentiment of the sentence is negative.
92.86
PromptSource
@@
Does the following sentence have a positive or negative sentiment?
93.03
EvoPrompt
@@
Examine the movie reviews and classify them as either positive or negative.
95.61
CR Manual Instruction
@@
Please perform Sentiment Classification task. Given the sentence, assign a sentiment label from [’negative’, ’positive’]. Return label only without any other text.
91.40
Natural Instruction
@@
In this task, you are given sentences from movie reviews. The task is to classify a sentence as "great" if the sentiment of the sentence is positive or as "terrible" if the sentiment of the sentence is negative.
90.90
EvoPrompt
@@
Analyze customer reviews and categorize each sentence as either ’positive’ or ’negative’.
91.75
MR Manual Instruction
@@
Please perform Sentiment Classification task. Given the sentence, assign a sentiment label from [’negative’, ’positive’]. Return label only without any other text.
88.75
Natural Instruction
@@
In this task, you are given sentences from movie reviews. The task is to classify a sentence as "great" if the sentiment of the sentence is positive or as "terrible" if the sentiment of the sentence is negative.
89.60
EvoPrompt
@@
Identify if a movie review is positive or negative by accurately categorizing each input-output pair into either ’positive’ or ’negative’.
91.35
SST-5 Manual Instruction
@@
Please perform Sentiment Classification task. Given the sentence, assign a sentiment label from [’terrible’, ’bad’, ’okay’, ’good’, ’great’]. Return label only without any other text.
42.90
Natural Instruction
@@
In this task, you are given sentences from movie reviews. Based on the given review, classify it to one of the five classes: (1) terrible, (2) bad, (3) okay, (4) good, and (5) great.
48.64
EvoPrompt
@@
Have your friend evaluate the movie they had just seen and provide a summary opinion (e.g. terrible, bad, okay, good, or great) to determine the sentiment of the movie review.
52.26
AG’s News Manual Instruction
@@
Please perform News Classification task. Given the news item, assign a label from [’World’, ’Sports’, ’Business’, ’Tech’]. Return label only without any other text.
70.63
Natural Instruction
@@
In this task, you are given a news article. Your task is to classify the article to one out of the four topics "World", "Sports", "Business", "Tech" if the article"s main topic is relevant to the world, sports, business, and technology, correspondingly. If you are not sure about the topic, choose the closest option.
48.89
PromptSource
@@
What label best describes this news article?
45.43
EvoPrompt
@@
Assess the entire concept of the news story and choose from the World, Sports, Business or Tech categories to categorize it into the correct category.
76.21
TREC Manual Instruction
@@
Please perform Question Classification task. Given the question, assign a label from [’Description’, ’Entity’, ’Expression’, ’Human’, ’Location’, ’Number’]. Return label only without any other text.
50.60
Natural Instruction
@@
You are given a question. You need to detect which category better describes the question. Answer with "Description", "Entity", "Expression", "Human", "Location", and "Number".
55.00
PromptSource
@@
Which category best describes the following question? Choose from the following list: Description, Entity, Abbreviation, Person, Quantity, Location.
36.20
EvoPrompt
@@
Recognize the inputs (explanations, entities, or humans) and provide the suitable outputs (numbers, descriptions, or entities) to answer the questions in a way that is understandable for non-native English speakers.
68.00
Subj Manual Instruction
@@
Please perform Subjectivity Classification task. Given the sentence, assign a label from [’subjective’, ’objective’]. Return label only without any other text.
49.75
Natural Instruction
@@
In this task, you are given sentences from reviews. The task is to classify a sentence as "subjective" if the opinion of the sentence is subjective or as "objective" if the opinion of the sentence is objective.
52.55
EvoPrompt
@@
Construct input-output pairs to demonstrate the subjectivity of reviews and opinions, distinguishing between objective and subjective input while producing examples of personal opinions and illustrations of subjective views, so it can illustrate the subjectivity of judgments and perspectives.
77.60
@@
Manual Instruction Alpaca-7b
How would you rephrase that in a few words?
35.92/11.16/31.67
GPT
How would you rephrase that in a few words?
43.95/17.11/39.09
EvoPrompt Alpaca-7b
Carefully examine the text or listen to the conversation to identify the key ideas, comprehend the main idea, and summarize the critical facts and ideas in the concise language without any unnecessary details or duplication.
39.86/14.24/36.09
GPT
Reduce the core by reading or listening carefully to identify the main ideas and key points, so readers can comprehend the important concepts and essential information.
46.49/19.49/41.96
@@
Manual Instruction Alpaca-7b
Simplify the text.
43.03
GPT-3.5
Simplify the text.
43.80
EvoPrompt Alpaca-7b
Rewrite the input text into simple English to make it easier to comprehend for non-native English speakers.
46.67
GPT-3.5
Rewrite the given sentence to make it more accessible and understandable for both native and non-native English speakers.
47.40
@@ Appendix D Future Works
01
hyperbaton
Order adjectives correctly in English sentences.
Verify the answer by splitting it into components and inspecting each part closely and logically, so we can progress thoughtfully and methodically as we break the task into pieces and explore each part systematically and rationally to reach our goal.
81.20
02
temporal_sequences
Answer questions about which times certain events could have occurred.
Start by breaking this conundrum into manageable chunks, carefully analyzing each component of this problem and thoroughly inspecting each aspect collaboratively, tackling it together progressively to ensure the correct answer and the desired outcome.
78.80
03
object_counting
Questions that involve enumerating objects and asking the model to count them.
Examine this logically and assess this methodically, so that we can obtain a precise result by thinking critically and dissecting this math task systematically.
87.60
04
disambiguation_qa
Clarify the meaning of sentences with ambiguous pronouns.
First, let us ponder and start off by taking our time, going step by step, and using our logic to approach this before we dive into the answer.
71.20
05
logical_deduction_three_objects
A logical deduction task which requires deducing the order of a sequence of objects.
Let’s approach it cautiously, examining it thoroughly and methodically, and then approach it incrementally towards a resolution.
94.40
05
logical_deduction_five_objects
A logical deduction task which requires deducing the order of a sequence of objects.
Split the problem into steps and thoughtfully progress through them to find the answer after the proof.
65.20
05
logical_deduction_seven_objects
A logical deduction task which requires deducing the order of a sequence of objects.
Let’s take a step-by-step approach to systematically dissect this math task.
54.40
@@
06
causal_judgement
Answer questions about causal attribution.
@@
At first, let’s handle things cautiously and resolve this by examining every detail and dealing with one problem at a time.
65.78
07
date_understanding
Infer the date from context.
@@
Be realistic and practical like a detective, and use evidence to solve the problem in a logical, step-by-step approach.
85.60
08
ruin_names
@@
Select the humorous edit that ’ruins’ the input movie or musical artist name.
Break down a math task into smaller sections and solve each one.
69.60
09
word_sorting
Sort a list of words.
@@
Analyze each part of the problem logically to solve it like a detective.
56.40
10
geometric_shapes
@@
Name geometric shapes from their SVG paths.
We’ll methodically work through this problem together.
64.00
11
movie_recommendation
@@
Recommend movies similar to the given list of movies.
Before exploring the answer,
86.00
12
salient_translation_error_detection
@@
Detect the type of error in an English translation of a German source sentence.
Break down the problem into individual steps in order to solve it.
62.80
13
formal_fallacies
@@
Distinguish deductively valid arguments from formal fallacies.
Let’s be realistic and evaluate the situation systematically, tackling it gradually.
56.00
14
penguins_in_a_table
Answer questions about a table of penguins and their attributes.
@@
Let’s start by taking a rational and organized approach, breaking it down into smaller parts and thinking it through logically, while being realistic and handling it carefully and methodically to ensure the right solution.
84.25
15
dyck_languages
Correctly close a Dyck-n word.
@@
Let’s be realistic and solve this challenge carefully and slowly, taking it slow to complete it correctly, so we can be realistic and cautiously reach the goal.
44.40
16
multistep_arithmetic_two
@@
Solve multi-step arithmetic problems.
Before we dive into the answer,
51.60
17
navigate
Given a series of navigation instructions, determine whether one would end up back at the starting point.
@@
Let’s logically work together to systematically solve this math problem one step at a time in unison.
94.20
18
reasoning_about_colored_objects
Answer extremely simple questions about the colors of objects on a surface.
@@
Using a detective’s mindset, break down each element of this mathematical reasoning challenge one step at a time and reason like a detective to uncover the solution.
88.00
19
boolean_expressions
Evaluate the result of a random Boolean expression.
@@
Let’s gradually unravel this mathematical challenge by methodically addressing it by examining each element and investigating each factor.
90.80
20
tracking_shuffled_objects_three_objects
@@
A task requiring determining the final positions of a set of objects given their initial positions and a description of a sequence of swaps.
Progress slowly and carefully through this mathematical reasoning challenge one step at a time.
69.20
20
tracking_shuffled_objects_five_objects
@@
A task requiring determining the final positions of a set of objects given their initial positions and a description of a sequence of swaps.
Using a logical, step-by-step approach, work through this task to find the correct answer.
81.20
20
tracking_shuffled_objects_seven_objects
@@
A task requiring determining the final positions of a set of objects given their initial positions and a description of a sequence of swaps.
Examine this issue logically and in detail, step-by-step, analyzing each part of the problem one at a time.
84.80
21
sports_understanding
@@
Determine whether an artificially constructed sentence relating to sports is plausible or not.
Break down the problem into steps and start solving it.
96.80
22
snarks
Determine which of two sentences is sarcastic.
@@
Break down and analyze each part of the problem in a step by step way to ensure the right answer is obtained.
77.53

### v3

# EvoPrompt: Connecting LLMs with Evolutionary Algorithms Yields Powerful Prompt Optimizers

@@ C.4 Analysis of Prompt
SST-2 Manual Instruction Please perform Sentiment Classification task. Given the sentence, assign a sentiment label from [’negative’, ’positive’]. Return label only without any other text. 93.68
@@
Natural Instruction In this task, you are given sentences from movie reviews. The task is to classify a sentence as "great" if the sentiment of the sentence is positive or as "terrible" if the sentiment of the sentence is negative. 92.86
@@
PromptSource Does the following sentence have a positive or negative sentiment? 93.03
@@
EvoPrompt Examine the movie reviews and classify them as either positive or negative. 95.61
@@
CR Manual Instruction Please perform Sentiment Classification task. Given the sentence, assign a sentiment label from [’negative’, ’positive’]. Return label only without any other text. 91.40
@@
Natural Instruction In this task, you are given sentences from movie reviews. The task is to classify a sentence as "great" if the sentiment of the sentence is positive or as "terrible" if the sentiment of the sentence is negative. 90.90
@@
EvoPrompt Analyze customer reviews and categorize each sentence as either ’positive’ or ’negative’. 91.75
@@
MR Manual Instruction Please perform Sentiment Classification task. Given the sentence, assign a sentiment label from [’negative’, ’positive’]. Return label only without any other text. 88.75
@@
Natural Instruction In this task, you are given sentences from movie reviews. The task is to classify a sentence as "great" if the sentiment of the sentence is positive or as "terrible" if the sentiment of the sentence is negative. 89.60
@@
EvoPrompt Identify if a movie review is positive or negative by accurately categorizing each input-output pair into either ’positive’ or ’negative’. 91.35
@@
SST-5 Manual Instruction Please perform Sentiment Classification task. Given the sentence, assign a sentiment label from [’terrible’, ’bad’, ’okay’, ’good’, ’great’]. Return label only without any other text. 42.90
@@
Natural Instruction In this task, you are given sentences from movie reviews. Based on the given review, classify it to one of the five classes: (1) terrible, (2) bad, (3) okay, (4) good, and (5) great. 48.64
@@
EvoPrompt Have your friend evaluate the movie they had just seen and provide a summary opinion (e.g. terrible, bad, okay, good, or great) to determine the sentiment of the movie review. 52.26
@@
AG’s News Manual Instruction Please perform News Classification task. Given the news item, assign a label from [’World’, ’Sports’, ’Business’, ’Tech’]. Return label only without any other text. 70.63
@@
Natural Instruction In this task, you are given a news article. Your task is to classify the article to one out of the four topics "World", "Sports", "Business", "Tech" if the article"s main topic is relevant to the world, sports, business, and technology, correspondingly. If you are not sure about the topic, choose the closest option. 48.89
@@
PromptSource What label best describes this news article? 45.43
@@
EvoPrompt Assess the entire concept of the news story and choose from the World, Sports, Business or Tech categories to categorize it into the correct category. 76.21
@@
TREC Manual Instruction Please perform Question Classification task. Given the question, assign a label from [’Description’, ’Entity’, ’Expression’, ’Human’, ’Location’, ’Number’]. Return label only without any other text. 50.60
@@
Natural Instruction You are given a question. You need to detect which category better describes the question. Answer with "Description", "Entity", "Expression", "Human", "Location", and "Number". 55.00
@@
PromptSource Which category best describes the following question? Choose from the following list: Description, Entity, Abbreviation, Person, Quantity, Location. 36.20
@@
EvoPrompt Recognize the inputs (explanations, entities, or humans) and provide the suitable outputs (numbers, descriptions, or entities) to answer the questions in a way that is understandable for non-native English speakers. 68.00
@@
Subj Manual Instruction Please perform Subjectivity Classification task. Given the sentence, assign a label from [’subjective’, ’objective’]. Return label only without any other text. 49.75
@@
Natural Instruction In this task, you are given sentences from reviews. The task is to classify a sentence as "subjective" if the opinion of the sentence is subjective or as "objective" if the opinion of the sentence is objective. 52.55
@@
EvoPrompt Construct input-output pairs to demonstrate the subjectivity of reviews and opinions, distinguishing between objective and subjective input while producing examples of personal opinions and illustrations of subjective views, so it can illustrate the subjectivity of judgments and perspectives. 77.60
@@
Manual Instruction Alpaca-7b How would you rephrase that in a few words? 35.92/11.16/31.67
GPT How would you rephrase that in a few words? 43.95/17.11/39.09
EvoPrompt Alpaca-7b Carefully examine the text or listen to the conversation to identify the key ideas, comprehend the main idea, and summarize the critical facts and ideas in the concise language without any unnecessary details or duplication. 39.86/14.24/36.09
GPT Reduce the core by reading or listening carefully to identify the main ideas and key points, so readers can comprehend the important concepts and essential information. 46.49/19.49/41.96
@@
Manual Instruction Alpaca-7b Simplify the text. 43.03
GPT-3.5 Simplify the text. 43.80
EvoPrompt Alpaca-7b Rewrite the input text into simple English to make it easier to comprehend for non-native English speakers. 46.67
GPT-3.5 Rewrite the given sentence to make it more accessible and understandable for both native and non-native English speakers. 47.40
@@ Appendix D Future Works
01 hyperbaton Order adjectives correctly in English sentences. Verify the answer by splitting it into components and inspecting each part closely and logically, so we can progress thoughtfully and methodically as we break the task into pieces and explore each part systematically and rationally to reach our goal. 81.20
02 temporal_sequences Answer questions about which times certain events could have occurred. Start by breaking this conundrum into manageable chunks, carefully analyzing each component of this problem and thoroughly inspecting each aspect collaboratively, tackling it together progressively to ensure the correct answer and the desired outcome. 78.80
03 object_counting Questions that involve enumerating objects and asking the model to count them. Examine this logically and assess this methodically, so that we can obtain a precise result by thinking critically and dissecting this math task systematically. 87.60
04 disambiguation_qa Clarify the meaning of sentences with ambiguous pronouns. First, let us ponder and start off by taking our time, going step by step, and using our logic to approach this before we dive into the answer. 71.20
05 logical_deduction_three_objects A logical deduction task which requires deducing the order of a sequence of objects. Let’s approach it cautiously, examining it thoroughly and methodically, and then approach it incrementally towards a resolution. 94.40
05 logical_deduction_five_objects A logical deduction task which requires deducing the order of a sequence of objects. Split the problem into steps and thoughtfully progress through them to find the answer after the proof. 65.20
05 logical_deduction_seven_objects A logical deduction task which requires deducing the order of a sequence of objects. Let’s take a step-by-step approach to systematically dissect this math task. 54.40
@@
06 causal_judgement Answer questions about causal attribution. At first, let’s handle things cautiously and resolve this by examining every detail and dealing with one problem at a time. 65.78
@@
07 date_understanding Infer the date from context. Be realistic and practical like a detective, and use evidence to solve the problem in a logical, step-by-step approach. 85.60
@@
08 ruin_names Select the humorous edit that ’ruins’ the input movie or musical artist name. Break down a math task into smaller sections and solve each one. 69.60
@@
09 word_sorting Sort a list of words. Analyze each part of the problem logically to solve it like a detective. 56.40
@@
10 geometric_shapes Name geometric shapes from their SVG paths. We’ll methodically work through this problem together. 64.00
@@
11 movie_recommendation Recommend movies similar to the given list of movies. Before exploring the answer, 86.00
@@
12 salient_translation_error_detection Detect the type of error in an English translation of a German source sentence. Break down the problem into individual steps in order to solve it. 62.80
@@
13 formal_fallacies Distinguish deductively valid arguments from formal fallacies. Let’s be realistic and evaluate the situation systematically, tackling it gradually. 56.00
@@
14 penguins_in_a_table Answer questions about a table of penguins and their attributes. Let’s start by taking a rational and organized approach, breaking it down into smaller parts and thinking it through logically, while being realistic and handling it carefully and methodically to ensure the right solution. 84.25
@@
15 dyck_languages Correctly close a Dyck-n word. Let’s be realistic and solve this challenge carefully and slowly, taking it slow to complete it correctly, so we can be realistic and cautiously reach the goal. 44.40
@@
16 multistep_arithmetic_two Solve multi-step arithmetic problems. Before we dive into the answer, 51.60
@@
17 navigate Given a series of navigation instructions, determine whether one would end up back at the starting point. Let’s logically work together to systematically solve this math problem one step at a time in unison. 94.20
@@
18 reasoning_about_colored_objects Answer extremely simple questions about the colors of objects on a surface. Using a detective’s mindset, break down each element of this mathematical reasoning challenge one step at a time and reason like a detective to uncover the solution. 88.00
@@
19 boolean_expressions Evaluate the result of a random Boolean expression. Let’s gradually unravel this mathematical challenge by methodically addressing it by examining each element and investigating each factor. 90.80
@@
20 tracking_shuffled_objects_three_objects A task requiring determining the final positions of a set of objects given their initial positions and a description of a sequence of swaps. Progress slowly and carefully through this mathematical reasoning challenge one step at a time. 69.20
@@
20 tracking_shuffled_objects_five_objects A task requiring determining the final positions of a set of objects given their initial positions and a description of a sequence of swaps. Using a logical, step-by-step approach, work through this task to find the correct answer. 81.20
@@
20 tracking_shuffled_objects_seven_objects A task requiring determining the final positions of a set of objects given their initial positions and a description of a sequence of swaps. Examine this issue logically and in detail, step-by-step, analyzing each part of the problem one at a time. 84.80
@@
21 sports_understanding Determine whether an artificially constructed sentence relating to sports is plausible or not. Break down the problem into steps and start solving it. 96.80
@@
22 snarks Determine which of two sentences is sarcastic. Break down and analyze each part of the problem in a step by step way to ensure the right answer is obtained. 77.53